### v1

# Simultaneous extension of the polynomial Central Sets Theorem and the polynomial Milliken-Taylor theorem

Abstract
In this article we find the polynomial extension of Beigelböck’s [1] multidimensional Central Sets Theorem (which was a common extension of Furstenberg’s Central Sets Theorem and Milliken-Taylor theorem). Then we derive a common extension of Bergelson, Hindman, and William’s [2] polynomial Milliken-Taylor theorem, and the polynomial multidimensional Central Sets Theorem. We also find a sufficient condition over two distinct multidimensional polynomials for which their Milliken-Taylor image lies in the same color. This result gives a partial polynomial extension of Deuber, Hindman, Leader, and Lefmann’s [5] result on separating the Milliken-Taylor system.
Mathematics subject classification 2020: 05D10, 05C55, 22A15, 54D35.
Keywords: Central sets, Central Sets Theorem, Milliken-Taylor system, Algebra of the Stone-Čech Compactifications of discrete semigroups.
1   Introduction
Throughout our article by $\mathbb{N}$, we denote the set of all positive integers. For any $r\in\mathbb{N},$ a $r$-coloring of a set $X$ is a partition of $X$ into $r$ disjoint sets. By finite coloring we mean any $r$-coloring, for any $r\in\mathbb{N}$. Let $(S,\cdot)$ be a semigroup, and $\mathcal{F}$ be a family of subsets of $S.$ We say $\mathcal{F}$ is partition regular (P.R.) if, for any finite coloring of $S,$ there exists a member of $F\in\mathcal{F}$ in the same color, i.e. $F$ is monochromatic. A cornerstone result in arithmetic Ramsey theory is the van der Waerden’s theorem [21], which states that the set of all arithmetic progression of a given length is P.R. Another important theorem was due to Schur [19] which states that the collection $\{\{x,y,x+y\}:x\neq y\in\mathbb{Z}\}$ is P.R. A nonlinear version of van der Waerden’s theorem, known as the Polynomial van der Waerden’s Theorem [^3] Actually, the authors proved a generalized version of the much stronger Szémeredi’s Theorem, but we are not going to discuss it in this paper. was established by Bergelson and Liebman in [3], using the methods of topological dynamics and PET induction. Throughout this article, we assume $\mathbb{P}$ be the set of all polynomials with no constant term.
Theorem 1.1 (Polynomial van der Waerden Theorem).
Let $r\in\mathbb{N},$ and $\mathbb{N}=\bigcup_{i=1}^{r}C_{i}$ be a $r$-coloring of $\mathbb{N}$. Then for any finite collection of polynomials $F$ in $\mathbb{P}$, there exist $a,d\in\mathbb{N}$ and $1\leq j\leq r$ such that $\{a+p(d):p\in F\}\subset C_{j}.$
An infinitary extension of Schur’s theorem inspires the following definition of IP sets. For any set $X$, let $\mathcal{P}_{f}\left(X\right)$ be the collection of all nonempty finite subsets of $X.$
Definition 1.2 (IP Sets).
Let $(S,+)$ be a commutative semigroup. A set $A\subseteq S$ is said to be an $IP$ set if there exists an injective sequence $\langle x_{n}\rangle_{n\in\mathbb{N}}$ in $S$ such that
$A=FS\left(\langle x_{n}\rangle_{n\in\mathbb{N}}\right)=\left\{\sum_{n\in\alpha}x_{n}:\alpha\in\mathcal{P}_{f}\left(\mathbb{N}\right)\right\}.$
Abbreviately for $\alpha\in\mathcal{P}_{f}\left(\mathbb{N}\right),$ we write $x_{\alpha}=\sum_{n\in\alpha}x_{n}$.
The following theorem is a cornerstone theorem in Ramsey theory, often known as the Hindman theorem [9].
Theorem 1.3 (Hindman Theorem).
If $(S,+)$ is a commutative semigroup, then for every finite coloring of $S$, there exists a monochromatic IP set.
For details on the classical approach to the Ramsey theory we refer to the book [8]. In our article we use Ultrafilter theory. So, before we proceed, we need to recall some basic results from the Stone-Čech compactification of discrete semigroups, sometimes known as Topological algebra/ theory of ultrafilter.
Topological algebra
Ultrafilters are set-theoretic objects which are intimately related to the Ramsey theory. In this section, we give a brief introduction to this theory. For details we refer to the book [13] to the readers.
A filter $\mathcal{F}$ over any nonempty set $X$ is a collection of subsets of $X$ such that
  1. 1.
    $\emptyset\notin\mathcal{F}$, and $X\in\mathcal{F}$,
  2. 2.
    $A\in\mathcal{F}$, and $A\subseteq B$ implies $B\in\mathcal{F},$
  3. 3.
    $A,B\in\mathcal{F}$ implies $A\cap B\in\mathcal{F}.$
Using Zorn’s lemma we can guarantee the existence of maximal filters which are called ultrafilters. Any ultrafilter $p$ has the following property:
  • if $X=\cup_{i=1}^{r}A_{i}$ is any finite partition of $X$, then there exists $i\in\{1,2,\ldots,r\}$ such that $A_{i}\in p.$
Let $(S,\cdot)$ be any discrete semigroup. Let $\beta S$ be the collection of all ultrafilters. For every $A\subseteq S,$ define $\overline{A}=\{p\in\beta S:A\in p\}.$ Now one can check that the collection $\{\overline{A}:A\subseteq S\}$ forms a basis for a topology. This basis generates a topology over $\beta S.$ We can extend the operation ``"``"`$\cdot$"` ` ⋅ " of $S$ over $\beta S$ as: for any $p,q\in\beta S,$ $A\in p\cdot q$ if and only if $\{x:x^{-1}A\in q\}\in p.$ With this operation ``"``"`$\cdot$"` ` ⋅ ", $(\beta S,\cdot)$ becomes a compact Hausdorff right topological semigroup. One can show that $\beta S$ is nothing but the Stone-Čech compactification of $S.$ Hence Ellis’s theorem guarantees that there exist idempotents in $(\beta S,\cdot)$. The set of all idempotents in $(\beta S,\cdot)$ is denoted by $E\left((\beta S,\cdot)\right).$ It can be shown that every member of the idempotents of $(\beta S,\cdot)$ contains an $IP$ set, which means every idempotent witnesses Hindman’s theorem. Using Zorn’s lemma one can show that $(\beta S,\cdot)$ contains minimal left ideals (minimal w.r.t. the inclusion). A well-known fact is that the union of such minimal left ideals is a minimal two-sided ideal, denoted by $K(\beta S,\cdot).$ Here we recall a few well-known classes of sets that are relevant to our work.
Definition 1.4.
Let $(S,\cdot)$ be a semigroup, let $n\in\mathbb{N}$ and let $A\subseteq S$. We say that
  • $A$ is a thick set if for any finite subset $F\subset S$, there exists an element $x\in S$ such that $Fx=\{fx:f\in F\}\subset A$;
  • $A$ is a syndetic set if there exists a finite set $F\subset S$ such that $S=\bigcup_{x\in F}x^{-1}A$, where $x^{-1}A=\{y:xy\in A\}$;
  • $A$ is piecewise syndetic set if there exists a finite set $F\subset S$ such that $\bigcup_{x\in F}x^{-1}A$ is a thick set. It is well known that $A$ is piecewise syndetic if and only if there exists $p\in K(\beta S,\cdot)$ such that $A\in p.$
  • $A$ is central set if it belongs to a minimal idempotent in $\beta S$.
It can be proved that a set $A$ is thick if and only if there exists a left ideal $L$ such that $L\subseteq\overline{A}$. And a set $A$ is syndetic if and only if for every left ideal $L$, $L\cap\overline{A}\neq\emptyset.$ Note that if $f:(S,\cdot)\rightarrow(T,\cdot)$ be any map between discrete semigroups. Then $\tilde{f}:\beta S\rightarrow\beta T$ be the continuous extension of $f$ defined by $\tilde{f}(p)=\tilde{f}(p-\lim x)=p-\lim f(x),$ where $p-\lim f(x)=y$ if and only if for every neighbourhood $U$ of $y$, $\{x:f(x)\in U\}\in p.$
Tensor product of ultrafilters
The tensor product of ultrafilters was introduced in [15]. Let $(S,\cdot)$ and $(T,\cdot)$ be two discrete semigroups. In this section, we recall materials from [14].
Definition 1.5.
Let $p\in(\beta S,\cdot)$ and $q\in(\beta T,\cdot)$ be two ultrafilters. Then the tensor product of $p$ and $q$ is defined as
$p\otimes q=\{A\subseteq S\times T:\{x\in S:\{y:(x,y)\in A\}\in q\}\in p\}$
where $s\in S$, $t\in T.$
Clearly $p\otimes q$ is a member of $\beta(S\times T).$ Tensor products can be characterized in terms of limits as follows.
$p\otimes q=\lim_{s\rightarrow p}\lim_{t\rightarrow q}(s,t).$
We can also extend the above definition of tensor products for multiple ultrafilters. The following definition uses inductive arguments.
Definition 1.6.
[2, Definition 1.15] Let $k\in\mathbb{N}$ and for $i\in\{1,2,\ldots,k\}$, let $S_{i}$ be a semigroup and let $p_{i}\in\beta S_{i}.$ We define $\bigotimes_{i=1}^{k}p_{i}\in\beta\left(\bigtimes_{i=1}^{k}S_{i}\right)$ as follows
  1. 1.
    $\bigotimes_{i=1}^{1}p_{i}=p_{1},$
  2. 2.
    given $k\in\mathbb{N}$, and $A\subseteq\bigtimes_{i=1}^{k+1}S_{i}$, $A\in\bigotimes_{i=1}^{k+1}p_{i}$ if and only if
    $\left\{(x_{1}\ldots,x_{k})\in\bigtimes_{i=1}^{k}S_{i}:\left\{x_{k+1}\in S_{k+1}:(x_{1}\ldots,x_{k+1})\in A\right\}\in p_{k+1}\right\}\in\bigotimes_{i=1}^{k}p_{i}.$
In a similar fashion we can say $\bigotimes_{i=1}^{k}p_{i}=\lim_{s_{1}\rightarrow p_{1}}\cdots\lim_{s_{k}\rightarrow p}(s_{1},\ldots,s_{k}).$
In the literature, the tensor product trick has been used to prove multi-dimensional Ramsey theoretic results including Milliken-Taylor theorem, and it’s polynomial extensions. For details, one can see [2, 13, 16]. In [1], Beigelböck proved a simultaneous extension of the Central Sets Theorem and Milliken-Taylor theorem using a little different approach described in [2, Chapter 18]. One of our main motivations is to study the tensor product of minimal idempotent ultrafilters, which was not extensively covered earlier. As a consequence we are able to find polynomial extensions of several multidimensional results in Ramsey theory. Before we proceed, we need the following technical Lemma, which is essential for our purpose.
Let $X$ be any nonempty set and let $\mathcal{P}(X)$ be the power set of $X.$
Lemma 1.7 (Lifting lemma).
[2, Lemma 4.3] Let $(S,\cdot)$ be a semigroup, let $m\in\mathbb{N}$, and let $p\in\beta S.$ Let $A\in\bigotimes_{j=1}^{m}p.$ Then for $j\in\{1,2,\ldots,m\}$ there exists $D_{j}:S^{j-1}\rightarrow\mathcal{P}(S)$ such that
  1. 1.
    for $j\in\{1,2,\ldots,m\},$ if for each $s\in\{1,2,\ldots,j-1\},$ $w_{s}\in D_{s}(w_{1},\ldots,w_{s-1}),$ then $D_{j}(w_{1},\ldots,w_{j-1})\in p;$
  2. 2.
    if for each $s\in\{1,2,\ldots,m\},$ $w_{s}\in D_{s}(w_{1},\ldots,w_{s-1}),$ then $(w_{1},\ldots,w_{m})\in A.$
Central Sets Theorem and its polynomial extensions
Before we proceed to our results, we need to recall the Central Sets Theorem. After the foundation of both van der Waerden’s and Hindman’s theorem, an immediate question appears if one can find a joint extension of both of these theorems. In [6], using the methods of Topological dynamics, Furstenberg defined the notions of Central Sets and proved that if $\mathbb{N}$ is finitely colored, then one of the color classes is Central. Then he proved a joint extension of both the van der Waerden’s and Hindman’s theorem, known as the Central Sets Theorem.
Theorem 1.8 (Central Sets Theorem).
Let $l\in\mathbb{N}$, and $A\subseteq\mathbb{N}$ be a central set. For each $i\in\{1,2,\ldots,l\}$ let $\langle x_{i,m}\rangle_{m=1}^{\infty}$ be a sequence in $\mathbb{\mathbb{Z}}$. Then there exist sequences $\langle b_{m}\rangle_{m=1}^{\infty}$ in $\mathbb{N}$ and $\langle K_{m}\rangle_{m=1}^{\infty}$ in $\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
  1. 1.
    For each $m$, $\max K_{m}<\min K_{m+1}$ and
  2. 2.
    For each $i\in\{1,2,\ldots,l\}$ and $H\in\mathcal{P}_{f}\left(\mathbb{N}\right)$, $\sum_{m\in H}(b_{m}+\sum_{t\in K_{m}}x_{i,t})\in A$.
Later the above theorem was generalized in several directions. For details about the developments and their combinatorial applications, we refer to the book [13]. Before we state the stronger version of Theorem 1.8, we need the following definitions from [4].
Definition 1.9.
($J$-set)
  1. 1.
    $A$ is called a $J$-set if and only if for every $H\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, there exists $a\in\mathbb{N}$ and $\beta\in\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that for all $f\in H,$ $a+\sum_{t\in\beta}f(t)\in A.$
  2. 2.
    $\mathcal{J}=\{p\in\beta\mathbb{N}:(\forall A\in p)\,A\text{ is a }J\text{ set}\}.$
It can be proved that $J$ sets are partition regular, and so $\mathcal{J}\neq\emptyset$. Again it is easy to verify that $(\mathcal{J},+)$ is a two sided ideal of $(\beta\mathbb{N},+)$, hence contains $K(\beta\mathbb{N},+).$ A set will be called $C$-set if it is a member of any idempotent in $(\mathcal{J},+)$. The following stronger version of the Central Sets Theorem was due to De, Hindman, and Strauss [4].
Theorem 1.10 (Stonger Central Sets Theorem, [4]).
Let $\tau=^{\mathbb{N}}\mathbb{N}$, and $C\subseteq\mathbb{N}$ be $C$ set. Then there exists functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
  1. 1.
    if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneqq G$ then $\max H(F)<\min H(G)$, and
  2. 2.
    whenever $m\in\mathbb{N}$, $G_{1},G_{2},\ldots,G_{m}\in\mathcal{P}_{f}(\tau)$, $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{m}$ and for each $i\in\{1,2,\ldots,m\}$, $f_{i}\in G_{i}$, one has
    $\sum_{i=1}^{m}\bigl{(}\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)\bigr{)}\in C.$
In fact it was also shown in [4] that if a set $A\subset\mathbb{N}$ contains the conclusion of Theorem 1.10, then it is a $C$ set.
Polynomial extension
In [11, 12] Hindman and McCutcheon extensively studied the algebraic structure of VIP systems which are the generalizations of usual polynomials. They proved several results including the polynomial extension of Theorem 1.8. This was the joint extension of the polynomial van der Waerden’s theorem and the central sets theorem.
Theorem 1.11 (Polynomial Central Sets Theorem).
Let $F\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, let $T\in\mathcal{P}_{f}(\mathbb{P})$ and let $A$ be a central subset of $\mathbb{N}$. Then there exist sequences $\langle b_{n}\rangle_{n=1}^{\infty}$ in $\mathbb{N}$ and $\langle H_{n}\rangle_{n=1}^{\infty}$ in $\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
  1. 1.
    for each $n\in\mathbb{N}$, $\max H_{n}<\min H_{n+1}$ and
  2. 2.
    for each $f\in F,$ each $P\in T$ and each $K\in\mathcal{P}_{f}\left(\mathbb{N}\right)$
$\sum_{n\in K}b_{n}+P\left(\sum_{n\in K}\sum_{t\in H_{n}}f(t)\right)\in A.$
Later in [7], both of the authors of this paper and Luperi Bagilini found the polynomial extension of Theorem 1.10. To state this result we need the following definition.
Definition 1.12.
($J_{p}$-set:)
  1. 1.
    $A$ is called a $J_{p}$-set if and only if for every $F\in\mathcal{P}_{f}(\mathbb{P})$, and every $H\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, there exists $a\in\mathbb{N}$ and $\beta\in\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that for all $P\in F$ and all $f\in H,$ $a+P\left(\sum_{t\in\beta}f(t)\right)\in A$
  2. 2.
    $\mathcal{J}_{p}=\{p\in\beta\mathbb{N}:(\forall A\in p)\,A\text{ is a }J_{p}\text{ set}\}.$
We don’t know whether $J_{p}$ sets are P.R., but it can be shown that every piecewise syndetic set is a $J_{p}$ set. Hence $\mathcal{J}_{p}\neq\emptyset$. Now one can show that $(\mathcal{J}_{p},+)$ is a two sided ideal of $(\beta\mathbb{N},+)$. We call a set a $C_{p}$ set if it belongs to some idempotents of $(\mathcal{J}_{p},+)$.
Theorem 1.13 (Stronger polynomial Central Sets Theorem).
[7, Theorem 11] Let $A$ be a $C_{p}$-set and let $T\in\mathcal{P}_{f}(\mathbb{P})$. There exist functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}(\mathbb{N})$ such that
  1. 1.
    if $G,K\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$ and $G\subsetneqq K$ then $\max H(G)<\min H(K)$ and
  2. 2.
    if $n\in\mathbb{N}$, $G_{1},G_{2},\cdots,G_{n}\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, $G_{1}\subsetneqq G_{2}\subsetneqq\cdots\subsetneqq G_{n}$ and for all $i\in\{1,2,\ldots,n\}$, $f_{i}\in G_{i}$, then for all $P\in T$,
    $\sum_{i=1}^{n}\alpha(G_{i})+P\left(\sum_{i=1}^{n}\sum_{t\in H(G_{i})}f_{i}(t)\right)\in A.$
However likewise in Theorem 1.10, the converse direction of Theorem 1.13 is not known. By observing that any $\beta\in\mathcal{P}_{f}\left(\mathbb{N}\right)$ is a subset of $\{1,\dots,n\}$ for some large enough $n\in\mathbb{N}$, we deduce the following seemingly stronger version of Theorem 1.13.
Corollary 1.14.
[7, Corollary 12] Let $A$ be a $C_{p}$-set and $T\in\mathcal{P}_{f}(\mathbb{P})$. Then there exist $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N},H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathcal{P}_{f}(\mathbb{N})$ such that
  1. 1.
    if $F,G\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N}),F\subset G$, then $\max H(F)<\min H(G)$;
  2. 2.
    if $\langle G_{n}\rangle_{n\in\mathbb{N}}$ is a sequence in $\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$ such that $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$ and $f_{i}\in G_{i},i\in\mathbb{N}$, then for all $P\in T,$
    $\sum_{i\in\beta}\alpha(G_{i})+P\left(\sum_{i\in\beta}\sum_{t\in H(G_{i})}f_{i}(t)\right)\in A.$
    for all $\beta\in\mathcal{P}_{f}\left(\mathbb{N}\right)$.
Milliken-Taylor Theorem and its extensions
In $1975$-$76$, Milliken and Taylor independently proved a joint extension of both Hindman theorem, and Ramsey’s theorem. To state the Milliken-Taylor Theorem, we need to introduce some notation. Given $F,G\in\mathcal{P}_{f}(\mathbb{N})$, we write $F<G$ to mean that $\max F<\min G.$ When we say that a sequence $\langle H_{n}\rangle_{n}$ in $\mathcal{P}_{f}(\mathbb{N})$ is increasing if for every $n\in\mathbb{N}$, $\max H_{n}<\min H_{n+1}.$ We present the different terminology because these special cases arise frequently.
Definition 1.15.
[2, Definition 1.3] Let $k\in\mathbb{N}$.
  1. 1.
    For any set $X,$ $[X]^{k}=\{A\subseteq X:|A|=k\}.$
  2. 2.
    For a sequence $\langle x_{n}\rangle_{n}$ in $\mathbb{N}$
    $[FS(\langle x_{n}\rangle_{n})]_{<}^{k}=\left\{\left(\sum_{t\in F_{1}}x_{t},\sum_{t\in F_{2}}x_{t},\ldots,\sum_{t\in F_{k}}x_{t}\right):F_{1}<F_{2}<\cdots<F_{k}\right\}.$
  3. 3.
    In a semigroup $(S,\cdot),$ $FP(\langle y_{n}\rangle_{n})$ is a product subsystem of $FP(\langle x_{n}\rangle_{n})$ if and only if there exists an increasing sequence $\langle H_{n}\rangle_{n}$ in $\mathcal{P}_{f}(\mathbb{N})$ such that for each $n\in\mathbb{N},$ $y_{n}=\Pi_{t\in H_{n}}x_{t}$ where the products $\Pi_{t\in H_{n}}x_{t}$ are computed in increasing order of indices. If the semigroup is commutative, we denote it by $(S,\cdot)$ and the product subsystem $FP(\langle y_{n}\rangle_{n})$ is denoted by $FS(\langle y_{n}\rangle_{n})$ referred by sum subsystem.
One of the fundamental theorems of Ramsey theory was due to Ramsey [18]:
Theorem 1.16 (Ramsey).
If $X$ is an infinite set, and $k\in\mathbb{N}$, then for every finite coloring of $[X]^{k}$, there exists an infinite set $Y\subseteq X$ such that the set $[Y]^{k}$ is monochromatic.
The following theorem was due to K. Milliken [17] and A. Taylor [20] generalizing both Ramsey’s theorem and Hindman’s theorem, known as the Milliken-Taylor theorem.
Theorem 1.17 (Milliken-Taylor).
Let $m,r\in\mathbb{N}$, and $[\mathbb{N}]^{m}=\bigcup_{i=1}^{r}C_{i}$ be a $r$-coloring. Let $\langle x_{n}\rangle_{n}$ be a sequence in $\mathbb{N}$. Then there exists $i\in\{1,2,\ldots,r\}$, and a sum subsystem $FS(\langle y_{n}\rangle_{n})$ of $FS(\langle x_{n}\rangle_{n})$ such that $[FS(\langle y_{n}\rangle_{n})]_{<}^{m}\subseteq C_{i}.$
To explain several extensions of the Milliken-Taylor theorem, we need the following notions of Milliken-Taylor system.
Definition 1.18 (Milliken-Taylor system).
Let $m\in\mathbb{N},$ let $\langle a_{j}\rangle_{j=1}^{m}$ and $\langle x_{n}\rangle_{n}$ be sequences in $\mathbb{N}.$ The Milliken-Taylor System determined by $\langle a_{j}\rangle_{j=1}^{m}$ and $\langle x_{n}\rangle_{n}$ is
$MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)=\left\{\sum_{j=1}^{m}a_{j}\sum_{t\in F_{j}}x_{t}:F_{1}<F_{2}<\cdots<F_{m}\right\}.$
The following theorem is another version of Theorem 1.17.
Theorem 1.19.
[2, Theorem 1.11.] Let $K\in\mathbb{N},$ let $\langle a_{j}\rangle_{j=1}^{k}$ and $\langle x_{n}\rangle_{n}$ be sequences in $\mathbb{N}.$ Let $g(z)=\sum_{j=1}^{k}a_{j}z$, and let $A\subseteq\mathbb{N}$. The following statements are equivalent.
  1. 1.
    There is an idempotent $p\in\bigcap_{m=1}^{\infty}\overline{FS\left(\langle x_{n}\rangle_{n=m}^{\infty}\right)}$ such that $A\in\tilde{g}(p)$
  2. 2.
    There is a sum subsystem $FS\left(\langle y_{n}\rangle_{n=1}^{\infty}\right)$ of $FS\left(\langle x_{n}\rangle_{n=1}^{\infty}\right)$ such that $MT\left(\langle a_{j}\rangle_{j=1}^{m},\langle x_{n}\rangle_{n}\right)\subseteq A.$
Using tensor product of ultrafilters one can find a seemingly stronger version of the Milliken-Taylor theorem.
Theorem 1.20.
[2, Theorem 1.17] Let $(S,\cdot)$ be a semigroup, and $m\in\mathbb{N},$ and let $A\subseteq\bigtimes_{i=1}^{m}S.$ The following statements are equivalent.
  1. 1.
    There is a sequence $\langle x_{n}\rangle_{n=1}^{\infty}$ in $S$ such that
    $\left\{\left(\sum_{t\in F_{1}}x_{t},\sum_{t\in F_{2}}x_{t},\ldots,\sum_{t\in F_{k}}x_{t}\right):F_{1}<F_{2}<\cdots<F_{k}\right\}\subseteq A.$
  2. 2.
    There is an idempotent $p\in\beta S$ such that $A\in\bigotimes_{i=1}^{m}p.$
In [1], Beigelböck simultaneously extends the combinatorial consequence of Theorem 1.8 and Theorem 1.20. Unfortunately, Beigelböck’s result does not address the reverse direction regarding ultrafilter characterization. The following theorem is his one.
Theorem 1.21.
Let $(S,\cdot)$ be a commutative semigroup and assume that there exists a nonprincipal minimal idempotent in $\beta S.$ For each $l\in\mathbb{N}$, let $\langle y_{l,n}\rangle_{n}$ be a sequence in $S.$ Let $k,r\in\mathbb{N}$ and let $[S]^{k}=\bigcup_{i}^{r}A_{i}.$ There exist $i\in\{1,2,\ldots,r\}$ a sequence $\langle a_{n}\rangle_{n}$ in $S$ and a sequence $\alpha_{0}<\alpha_{1}<\cdots<\alpha_{n}<\cdots$ in $\mathcal{P}_{f}(\mathbb{N})$ such that $\Big{[}FP\big{(}\langle a_{n}\prod_{t\in\alpha_{n}}y_{g(n),t}\rangle_{n}\big{)}\Big{]}_{<}^{k}\subset A_{i}$
Our first result is to improve Theorem 1.21. We simultaneously extend Theorem 1.10 and Theorem 1.26. In our article, we restrict ourselves to the set of natural numbers, but this technique can be mimicked for commutative semigroups.
Theorem 1.22.
Let $m\in\mathbb{N}$, and $A\subseteq\mathbb{N}$. Then the following two statements are equivalent.
  1. 1.
    there exists functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
    1. (a)
      if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneq G$ then $\max H(F)<\min H(G)$, and
    2. (b)
      if $n(\geq m)\in\mathbb{N}$, $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{m}$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m}\leq\{n\},$ in $\mathcal{P}_{f}(\mathbb{N})$ we have
      $\left\{\left(\sum_{i\in\beta_{1}}\left(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)\right),\cdots,\sum_{i\in\beta_{m}}\left(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)\right):f_{i}\in G_{i}\text{ for every }i\in\mathbb{N}\right)\right\}\subset A.$
  2. 2.
    There exists a idempotent $p$ in $(\mathcal{J},+)$, such that $A\in\bigotimes_{i=1}^{m}p.$
Similar to the Corollary 1.14, the following corollary of the above theorem gives an infinitary version.
Corollary 1.23.
Let $m\in\mathbb{N}$, and $A\subseteq\mathbb{N}$. Then the following two statements are equivalent.
  1. 1.
    there exists functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
    1. (a)
      if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneq G$ then $\max H(F)<\min H(G)$, and
    2. (b)
      whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$ we have
      $\left\{\left(\sum_{i\in\beta_{1}}\left(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)\right),\cdots,\sum_{i\in\beta_{m}}\left(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)\right):f_{i}\in G_{i}\text{ for every }i\in\mathbb{N}\right)\right\}\subset A.$
  2. 2.
    There exists a idempotent $p$ in $(\mathcal{J},+)$, such that $A\in\bigotimes_{i=1}^{m}p.$
Our next result polynomialize the Theorem 1.22. Unfortunately, we are unable to find the ultrafilter characterization of the next theorem. The main problem is that we don’t know any ultrafilter characterization of the polynomial Central Sets Theorem.
Theorem 1.24.
Let $m\in\mathbb{N}$, and let $p$ be a minimal idempotent in $(\beta\mathbb{N},+)$. Then for every $A\in\bigotimes_{i=1}^{m}p$, and $T\in\mathcal{P}_{f}(P)$, we have two functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$, and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathcal{P}_{f}(\mathbb{N})$ such that
  1. 1.
    if $G,K\in\mathcal{P}_{f}((^{\mathbb{N}}\mathbb{N})$, and $G\subsetneq K$, then $H(G)<H(K)$, and
  2. 2.
    if $n(\geq m)\in\mathbb{N}$, $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{m}$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m}\leq\{n\},$ in $\mathcal{P}_{f}(\mathbb{N})$, $f_{i}\in G_{i}$, and $P\in T$ we have
    $\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)\in A.$
In fact we may consider $p\in E\left(\mathcal{J}_{p},+\right).$
Again similar to the Corollary 1.14, the following corollary of the above theorem gives an infinitary version.
Corollary 1.25.
Let $m\in\mathbb{N}$, and let $p$ be a minimal idempotent in $(\beta\mathbb{N},+)$. Then for every $A\in\bigotimes_{i=1}^{m}p$, and $T\in\mathcal{P}_{f}(P)$, we have two functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$, and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathcal{P}_{f}(\mathbb{N})$ such that
  1. 1.
    if $G,K\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, and $G\subsetneq K$, then $H(G)<H(K)$, and
  2. 2.
    if $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{n}\subsetneq\cdots$ in $P_{f}(^{\mathbb{N}}\mathbb{N})$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m},$ in $\mathcal{P}_{f}(\mathbb{N})$, $f_{i}\in G_{i}$, and $P\in T$ we have
    $\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)\in A.$
In fact we may consider $p\in E\left(\mathcal{J}_{p},+\right).$
Polynomial extension of the Milliken-Taylor theorem
Let $m\in\mathbb{N}$, and let $h:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be any polynomial having no constant term. Then from [2, Theorem 3.2] it is clear that for every $p\in\beta\mathbb{N}$, $\tilde{h}\left(\bigotimes_{i=1}^{m}p\right)=h(p,\ldots,p).$ The following theorem is a corollary of [2, Theorem 3.2], and is the polynomial extension of Theorem 1.26.
Theorem 1.26.
[2, Theorem 1.11.] Let $m\in\mathbb{N},$ $h:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be any polynomial having no constant term, and let $A\subseteq\bigtimes_{i=1}^{m}\mathbb{N}.$ Then the following statements are equivalent.
  1. 1.
    There is a sequence $\langle x_{n}\rangle_{n=1}^{\infty}$ in $\mathbb{N}$ such that
    $\left\{h\left(\sum_{t\in F_{1}}x_{t},\sum_{t\in F_{2}}x_{t},\ldots,\sum_{t\in F_{k}}x_{t}\right):F_{1}<F_{2}<\cdots<F_{k}\right\}\subseteq A.$
  2. 2.
    There is an idempotent $p\in\beta S$ such that $A\in\tilde{h}\left(\bigotimes_{i=1}^{m}p\right).$
The following corollary immediately follows from [2, Theorem 3.2], and Corollary 1.23, and can be thought of as the polynomial extension of Corollary 1.23.
Corollary 1.27.
Let $m\in\mathbb{N},$ $h:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be any polynomial having no constant term, and let $A\subseteq\bigtimes_{i=1}^{m}\mathbb{N}.$ Then the following statements are equivalent.
  1. 1.
    there exists functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
    1. (a)
      if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneq G$ then $\max H(F)<\min H(G)$, and
    2. (b)
      whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$ we have
      $\left\{h\left(\sum_{i\in\beta_{1}}\left(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)\right),\cdots,\sum_{i\in\beta_{m}}\left(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)\right):f_{i}\in G_{i}\text{ for every }i\in\mathbb{N}\right)\right\}\in A.$
  2. 2.
    There exists a idempotent $p$ in $(\mathcal{J},+)$, such that $A\in\tilde{h}\left(\bigotimes_{i=1}^{m}p\right).$
In the above theorem, we did not consider any polynomial inside $h.$ This inspires us the following generalization of the Corollary 1.25.
Corollary 1.28.
Let $m\in\mathbb{N}$, $h:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be any polynomial having no constant term, and let $p$ be a minimal idempotent in $(\beta\mathbb{N},+)$. Then for every $A\in\tilde{h}\left(\bigotimes_{i=1}^{m}p\right)$, and $T\in\mathcal{P}_{f}(P)$, we have two functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$, and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathcal{P}_{f}(\mathbb{N})$ such that
  1. 1.
    if $G,K\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, and $G\subsetneq K$, then $H(G)<H(K)$, and
  2. 2.
    if $G_{1}\subsetneq G_{2}\subsetneq\dots\subsetneq G_{n}\subsetneq\cdots$ in $P_{f}(^{\mathbb{N}}\mathbb{N})$, then for each $\beta_{1}<\beta_{2}<\dots<\beta_{m},$ in $\mathcal{P}_{f}(\mathbb{N})$, $f_{i}\in G_{i}$, and $P\in T$ we have
    $h\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)\in A.$
In fact we may consider $p\in E\left(\mathcal{J}_{p},+\right).$
Proof.
Immediately follows from [2, Theorem 3.2], and Corollary 1.25. ∎
Structure of the paper:
In the next section, we prove Theorem 1.22. Let $i:\mathbb{N}\rightarrow\mathbb{N}$ be a polynomial defined by $i(x)=x$ for all $x\in\mathbb{N}.$ Then if we choose $T=\{i\}$ in Theorem 1.24, then a similar proof of Theorem 1.24 implies $(1)\implies(2)$ in Theorem 1.22. Hence we need to prove the reverse direction. So our first focus will be to prove $(2)\implies(1)$ in Theorem 1.22. Then we prove Theorem 1.24. Then we prove another version of Theorem 1.24. Then in the last section we address the polynomial extension of separating the Milliken-Taylor theorem, originally proved in [5].
2   Our proofs
As we promised before, in this section we will prove Theorem 1.22, and Theorem 1.24. Then we draw several new results using these theorems.
Proof of Theorem 1.22.
As mentioned earlier we are going to proof only $(2)\implies(1).$
Let $\alpha:\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{N}$ and $H:\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})\rightarrow\mathbb{P}_{f}(\mathbb{N})$ be two functions such that $\max H(F)<\min H(G)$ provided $F,G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$ and $F\subsetneq G$. For $G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$, let $T_{G}=\{\sum^{m}_{i=1}\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t):n\in\mathbb{N},\text{ each }G_{i}\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N}),G\subsetneq G_{1}\subsetneq G_{2}\cdots\subsetneq G_{n},\text{ and for each }i\in\{1,2,\ldots n\},f_{i}\in G_{i}\}$.
Now for $\alpha:\mathbb{P}_{f}(\mathbb{N}_{\mathbb{N}})\rightarrow\mathbb{N}$ and $H:\mathbb{P}_{f}(\mathbb{N})\rightarrow\mathbb{P}_{f}(\mathbb{N})$ with $\max H(F)<\min H(G)$ provided $F,G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$ and $F\subsetneq G$; define
$Q(\alpha,H)=\bigcap_{G\in\mathbb{P}_{f}(\mathbb{N}_{\mathbb{N}})}\overline{T}_{G}.$
Then $Q(\alpha,H)\neq\emptyset$ and is a subgroup of $(\beta\mathbb{N},+)$. Furthermore,
$K(Q(\alpha,H))\subseteq(\mathcal{J},+).$
We show, by downward induction on $l\in\{1,2,\ldots,m\}$ that for each $G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$,
$\left\{\sum_{i\in\beta_{l}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\sum_{i\in\beta_{l+1}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\ldots,\sum_{i\in\beta_{m}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)):\right.\\ \left.f_{i}\in G_{i},G\subseteq G_{i}\text{ for every }i\in\mathbb{N}\text{ and }\beta_{l}<\beta_{l+1}<\cdots<\beta_{m}\right\}\in\bigotimes_{i=l}^{m}p.$
For $l=m$, we have $T_{G}\in p$ so that,
$\left\{\sum_{i\in\beta_{m}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)):f_{i}\in G_{i},G\subseteq G_{i}\text{ for every }i\in\mathbb{N}\right\}\subset\bigotimes_{i=m}^{m}p.$
Therefore, let $l\in\{1,2,\ldots,m-1\}$ and assume that the statement is true for $l+1$. Now let, $G\in\mathbb{P}_{f}(^{\mathbb{N}}\mathbb{N})$ and
$A=\left\{\sum_{i\in\beta_{l}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\sum_{i\in\beta_{l+1}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\ldots,\sum_{i\in\beta_{m}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)):\right.\\ \left.f_{i}\in G_{i},G\subseteq G_{i}\text{ for every }i\in\mathbb{N}\text{ and }\beta_{l}<\beta_{l+1}<\cdots<\beta_{m}\right\}.$
We claim that,
$T_{G}\subseteq\{x\in\mathbb{N}:\{(x_{l+1},x_{l+2},\ldots,x_{m})\in\mathbb{N}^{m-l}\}:(x,x_{l+1},x_{l+2},\ldots,x_{m})\in A\}\in\bigotimes_{i=l+1}^{m}p_{i}.$
To this end let $b\in T_{G}$. Pick $G_{l}\in\mathbb{P}_{f}(\mathbb{N}_{\mathbb{N}})$ with $G\subseteq G_{l}$ such that $b=\sum_{i\in\beta_{l}}\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)$. Let $F=\cup_{i\in\beta_{l}}G_{i}$. Then
$\left\{\sum_{i\in\beta_{l+1}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\sum_{i\in\beta_{l+2}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)),\ldots,\sum_{i\in\beta_{m}}(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)):\right.\\ \left.f_{i}\in G_{i},F\subseteq G_{i}\text{ for every }i\in\mathbb{N}\text{ and }\beta_{l+1}<\beta_{l+2}<\cdots<\beta_{m}\right\}\subseteq-b+A.$
So, $\{(x_{l+1},x_{l+2},\ldots,x_{m})\in\mathbb{N}^{m-l}:(b,x_{l+1},x_{l+2},\ldots,x_{m})\in A\}\in\bigotimes_{i=l+1}^{m}p_{i}$ as required.
Now we will prove Theorem 1.24, but before that let us recall one fact about ultrafilters from [13]. We will apply this result. For any discrete semigroup $(S,\cdot)$, and $p\in E\in\left(\beta S,\cdot\right)$, if $A\in p=p\cdot p$, then from [13, Lemma 4.14], the set $A^{\star}=\{x\in A:-x+A\in p\}\in p.$
Proof of Theorem 1.24:.
Let $p\in E(K(\beta\mathbb{N},+))$, and $q=p\otimes p\otimes\dots\otimes p$, and let $A\in q$. By using Lifting lemma 1.7, pick for each $j\in\{1,\ldots,m\}$, some $D_{j}:\mathbb{N}^{j-1}\rightarrow\mathcal{P}(\mathbb{N})$ such that
  1. ($1$)
    for each $j\in\{1,2,\ldots,m\}$, and for each $s\in\{1,2,\ldots,j-1\},w_{s}\in D_{s}(w_{1},w_{2},\dots,w_{s-1})$, and then $D_{j}(w_{1},\dots,w_{j-1})\in p$, and
  2. ($2$)
    if for each $s\in\{1,2,\ldots,m\},w_{s}\in D_{s}(w_{1},w_{2},\ldots,w_{s-1})$, then $(w_{1},w_{2},\dots,w_{m})\in A$.
Now we proceed using the induction argument.
For the base case, let $K=\{f\}$. As $D_{1}(\emptyset)^{\star}\in P$, there exists $a\in\mathbb{N}$, and $\beta\in\mathcal{P}_{f}(\mathbb{N})$ such that for all $P\in T,a+P\left(\sum_{t\in\beta}f(t)\right)\in D_{1}(\emptyset)^{\star}$. Define $\alpha(K)=a$, and $H(K)=\beta$. Now assume that $|K|>1$, and that $\alpha(G)$, and $H(G)$ have been defined for all non-empty proper subsets $G$ of $K$ satisfying
  1. ($1$)
    If $G,K\in\mathcal{P}_{f}(\mathbb{N}^{\mathbb{N}})$ and $G\subsetneq K$, then $H(G)<H(K)$.
  2. ($2$)
    If $n\in\mathbb{N}$, $G_{1}\subsetneq G_{2}\subsetneq\ldots\subsetneq G_{n}\subsetneq K$, then for each $j\in\{1,2,\ldots,m\}$, and $\beta_{1}<\beta_{2}<\dots<\beta_{j}\leq\{n\}$,
    $\sum_{i\in\beta_{j}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{j}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\in$
    $D_{j}\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\dots,\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)^{\star}$
Let, $R=\left\{\sum_{i\in\beta_{j}}\sum_{t\in H(G_{i})}f_{i}(t):\beta_{j}\leq\{n\},f_{i}\in G_{i}\right\}$, $F=\bigcup\{H(G):\emptyset\neq G\subsetneq K\}$, and let $m=\max F$ and, let for each $j\in\{1,2,\ldots,m\}$,
$M_{j}=\left\{\sum_{i\in\beta_{j}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{j}}\sum_{t\in H(G_{i})}f_{i}(t)\right):\emptyset\neq G_{1}\subsetneq\ldots\subsetneq G_{n}\subsetneq K,P\in T,\text{ and for each }i\in\{1,\ldots,n\};f_{i}\in G_{i}\right\}.$
Then $M_{j}\subseteq D_{j}^{\star}$.
Let $P\in T,d\in R$, define the polynomial $Q_{p,d}$ as $Q_{p,d}(y)=P(y+d)-P(d)$. Clearly $Q_{p,d}\in\mathbb{P}$. Let $S=T\cup\{Q_{p,d}:P\in T\text{ and }d\in R\}$.
Define
$E=\bigcap^{m}_{j=1}\left\{D_{j}\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}\right)\right)^{\star}:\right.\\ \left.\emptyset\neq G_{1}\subsetneq\ldots\subsetneq G_{n}\subsetneq K,P\in T,\text{ and for each }i\in\{1,\dots,n\},f_{i}\in G_{i}\right\}$
$G=\bigcap_{\underset{j=1,\dots,m}{x\in M;}}-x+D_{j}\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\dots,\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)^{\star}\in P$
Now by [7, Lemma 10], pick $a\in\mathbb{N}$, and $\beta\in\mathcal{P}_{f}(\mathbb{N})$ such that $\min\beta>m$, and for all $Q\in S$, and all $f\in K$,
$a+Q\left(\sum_{t\in\beta}f(t)\right)\in G.$
Let $\alpha(K)=a$, and $H(K)=\beta$. Since $\min\beta>m$, $(1)$ is satisfied. To verify ($2$), let $n\in\mathbb{N}$, and let $G_{1},G_{2},\ldots,G_{n}\in\mathcal{P}_{f}(\mathbb{N}_{\mathbb{N}})$ with $G_{1}\subsetneq G_{2}\subsetneq\ldots\subsetneq G_{n}=K$, for each $i\in{1,\ldots,n}$, let $f_{i}\in G_{i}$, and let $P\in T$. If $j=1$, and $\beta_{1}<\{n\}$, then this follows from induction hypothesis. If $j=1$, and $\beta_{1}=\{n\}$, then $G_{n}=K$.
Now for all $P\in T,\alpha(K)+P\left(\sum_{t\in H(k)}f_{i}(t)\right)\in G\subset D_{1}^{\star}(\emptyset)$. Otherwise if, $\beta_{1}=\gamma\cup\{n\}$, then we can rename $\gamma$ as $\beta_{1}$, and $\{n\}$ as $\beta_{2}$. Now we can proceed as follows. Choose any $\beta_{1}<\beta_{2}<\dots<\beta_{j}<\{n\}$, where $2\leq j\leq m$. Then
$\sum_{i\in\beta_{j}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{j}}\sum_{i\in H(G_{i})}f_{i}(t)\right)=a+\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)+\\ P\left(\sum_{t\in\beta}f_{n}(t)+\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)-P\left(\sum_{t\in\beta}f_{n}(t)+\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\\ =a+x+Q_{p,d}\left(\sum_{t\in\beta}f_{n}(t)\right)\in D_{j}\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{t\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{j-1}}\alpha(G_{i})+\right.\\ \left.P\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right),$
where, $x=\sum_{i\in\beta_{j-1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)\right),d=\sum_{i\in\beta_{j-1}}\sum_{t\in H(G_{i})}f_{i}(t)$. ∎
The following theorem shows that the combinatorial consequence of our Corollary 1.25 is equivalent to a version of the polynomial extension of [13, Theorem 17.31]. The proof uses simple color induction arguments.
Theorem 2.1.
Let $T\in\mathcal{P}_{f}(\mathbb{P})$. Then the following two versions are equivalent.
  1. 1.
    Let $r,m\in\mathbb{N}$, and let $[\mathbb{N}]^{m}=\bigcup_{i=1}^{r}C_{i}$. Then there exists functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
    1. (a)
      if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneq G$ then $\max H(F)<\min H(G)$, and
    2. (b)
      whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $P\in T,$ $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$, we have
      $\left\{\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right):f_{i}\in G_{i}\text{ for every }i\in\mathbb{N}\right\}\subset C_{j}$
      for some $j\in\{1,2,\ldots,r\}.$
  2. 2.
    Let $r,m\in\mathbb{N}$, and let $a_{1},\dots,a_{m}\in\mathbb{N}$. Then for any $\mathbb{N}=\bigcup^{r}_{i=1}D_{i}$, there exists functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that for every $P\in T,$
    1. (a)
      if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneq G$ then $\max H(F)<\min H(G)$, and
    2. (b)
      whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$ we have for every $f_{i}\in G_{i},$ and $P\in T,$
      $a_{1}\cdot\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right)\right)+\cdots+a_{m}\cdot\left(\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(a_{i})}f_{i}(t)\right)\right)\in D_{j}$
      for some $j\in\{1,2,\ldots,r\}.$
Proof.
$(1)\implies(2)$ Let $\mathbb{N}=\bigcup^{r}_{i=1}D_{i}$, and induce a coloring $\left[\mathbb{N}\right]^{(m)}=\bigcup^{r}_{i=1}C_{i}$ by defining $(z_{1},z_{2},\dots,z_{m})\in C_{i}\Leftrightarrow a_{1}z_{1}+\dots+a_{m}z_{m}\in D_{i}$. Now (2) directly follows from (1).
$(2)\implies(1)$ Proof is similar. ∎
In the following theorem, we explicitly deduce some ultrafilters witnessing the polynomial extension of [13, Theorem 17.31]. However, we don’t know if the converse part of this theorem is true or not.
Theorem 2.2.
Let $m\in\mathbb{N},$ $a_{1},\dots,a_{m}\in\mathbb{N}$, $\,p\in E\left(K(\beta\mathbb{N},+)\right)$, and $A\in a_{1}p+\dots+a_{m}p$. Then for every $T\in\mathcal{P}_{f}(\mathbb{P})$, there exists functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
  1. 1.
    if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneq G$ then $\max H(F)<\min H(G)$, and
  2. 2.
    whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $P\in T,$ $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$, we have
    $a_{1}\cdot\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right)\right)+\cdots+a_{m}\cdot\left(\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(a_{i})}f_{i}(t)\right)\right)\in A.$
In fact we may consider $p\in E\left(\mathcal{J}_{p},+\right).$
Proof.
Let $\Phi:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be the map defined by $\Phi(x_{1},x_{2},\dots,x_{m})=\sum^{m}_{i=1}a_{i}x_{i}$. Clearly $\Phi$ is a continuous map, and so we may apply [13, Theorem 3.49]. Let $\widetilde{\Phi}:\beta(\mathbb{N}^{m})\rightarrow\beta\mathbb{N}$ be the continuous extension of $\Phi$. Let $p\in E(K(\beta\mathbb{N},+))$, and $\bigotimes_{i=1}^{m}p\in\beta(\mathbb{N}^{m})$. Now,
$\widetilde{\Phi}\left[\bigotimes_{i=1}^{m}\right]=\widetilde{\Phi}\left[p-\lim_{x_{1}}\cdots p-\lim_{x_{m}}(x_{1},\dots,x_{m})\right]$
$=p-\lim_{x_{1}}\cdots p-\lim_{x_{m}}\Phi(x_{1},\dots,x_{m})$
$=p-\lim_{x_{1}}\dots p-\lim_{x_{m}}\sum^{m}_{i=1}a_{i}x_{i}$
$=a_{1}\cdot p+\dots+a_{m}\cdot p.$
Let $A\in a_{1}p+\dots+a_{m}p=\widetilde{\Phi}\left[\bigotimes_{i=1}^{m}p\right]$. Hence there exists $B\in\bigotimes_{i=1}^{m}p$ such that $\Phi[B]\subseteq A$. Now from Corollary 1.25, our result follows.
The following corollary is an immediate consequence of Theorem 2.2.
Corollary 2.3.
Let $m\in\mathbb{N},$ $h:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be any polynomial with no constant term. Let $a_{1},\dots,a_{m}\in\mathbb{N}$, $\,p\in E\left(K(\beta\mathbb{N},+)\right)$, and $A\in\tilde{h}(a_{1}p+\dots+a_{m}p)$. Then for every $T\in\mathcal{P}_{f}(\mathbb{P})$, there exists functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
  1. 1.
    if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneq G$ then $\max H(F)<\min H(G)$, and
  2. 2.
    whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $P\in T,$ $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$, we have
    $h\left(a_{1}\cdot\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right)\right)+\cdots+a_{m}\cdot\left(\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(a_{i})}f_{i}(t)\right)\right)\right)\in A.$
In fact we may consider $p\in E\left(\mathcal{J}_{p},+\right).$
Proof.
Let $A\in\tilde{h}(a_{1}p+\dots+a_{m}p)$, and let $B\in a_{1}p+\dots+a_{m}p$ be such that $h[B]\subseteq A.$ Now Theorem 2.2 directly implies our result. ∎
3   Separation of polynomial Milliken-Taylor systems
In [5, 10], authors extensively studied and characterized the compatibility of two Milliken-Taylor systems over the set of integers. That means they addressed when two Milliken-Taylor systems generated by two distinct finite sets lie in the same color after a finite coloring of $\mathbb{N}.$ To address their result of [5], we need the notions of compressed sequence.
Definition 3.1.
  1. 1.
    Let $S=\{\overrightarrow{a}:\overrightarrow{a}\text{ is a finite sequence in }\mathbb{N}\}.$
  2. 2.
    The function $c:S\rightarrow S$ deletes any consecutive repeated terms.
  3. 3.
    An element $\overrightarrow{a}\in S$ is compressed if and only if $\overrightarrow{a}=c(\overrightarrow{a}).$
  4. 4.
    An equivalence relation on $S$ is defined by $a\approx b$ if and only if there exists a rational $\alpha$ such that $\alpha\cdot c(\overrightarrow{a})=c(\overrightarrow{b}).$
The following theorem gives a characterization of two distinct Milliken-Taylor systems lie in the same partition.
Theorem 3.2.
[5, Theorem 3.2,3.3] Let $\overrightarrow{a},\overrightarrow{b}\in S.$ Then the following are equivalent.
  1. 1.
    For every finite coloring there exists two sequences $\langle x_{n}\rangle_{n}$ and $\langle y_{n}\rangle_{n}$ such that $MT\left(\overrightarrow{a},\langle x_{n}\rangle_{n}\right)\bigcup MT\left(\overrightarrow{b},\langle y_{n}\rangle_{n}\right)$ is monochromatic.
  2. 2.
    $a\approx b.$
In [10], authors extended Theorem 3.2 over the set of integers $\mathbb{Z}.$ They proved Theorem 3.2 under the condition that $\overrightarrow{a},\overrightarrow{b}$ are two finite sets from $\mathbb{Z}\setminus\{0\}.$ From [2] an immediate question appear if there exists any polynomial extension of Theorem 3.2. Throughout this article we have seen the ultrafilter characterizations of polynomial patterns are rich in today’s technique. But here we prove the polynomial extension of one side of Theorem 3.2. Before that, we introduce a polynomial version of an equivalent compressed sequence.
Definition 3.3.
Let $m\in\mathbb{N}$, and $f,g:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be two polynomials of finite degree defined by $f(x_{1},\dots,x_{m})=\sum a_{i_{1}\cdots i_{dn}}x^{i_{1}}_{1}\cdots x^{i_{m}}_{m}$, and $g(x_{1},\dots,x_{m})=\sum b_{j_{1}\cdots j_{m}}x^{j_{1}}_{1}\cdots x^{j_{m}}_{m}$. We say $f$ and $g$ are equivalent, i.e. $f\approx g$ if and only if, there exists a rational $\alpha\in\mathbb{Q}$ such that,
$b_{i_{1}\cdots i_{m}}=\alpha^{i_{1}+\dots+i_{m}}a_{i_{1}\dots i_{m}}\forall\,(i_{1},\dots,i_{m}).$
For example the polynomials $f(x,y,z)=x^{3}+2xy^{2}z+xz+z^{2}$, and $g(x,y,z)=5x^{3}+10xy^{2}z+5xz+5z^{2}$ are equivalent.
Now we introduce the notion of the polynomial Milliken-Taylor system. This notion generalizes the notion of the Milliken-Taylor system 1.18.
Definition 3.4 (Polynomial Milliken-Taylor system).
Let $f:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be any polynomial with no constant term. Let $\langle x_{n}\rangle^{\infty}_{n=1}$ be any sequence in $\mathbb{N}$. Define
$MT\left(f,\langle x_{n}\rangle^{\infty}_{n=1}\right)=\left\{f\left(\sum_{i\in\beta_{1}}x_{i},\dots,\sum_{i\in\beta_{m}}x_{i}\right):\beta_{1}<\dots<\beta_{m}\right\}.$
The following theorem shows that if two polynomials are equivalent then their Milliken-Taylor image for some sequences lies in the same color after partitioning $\mathbb{N}$ into finitely many pieces.
Theorem 3.5 (Separating Polynomial Milliken-Taylor).
Let $r,m\in\mathbb{N}$, and $f,g:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be two finite degree polynomials with no constant term such that $f\approx g$. Then if $\mathbb{N}=\bigcup^{r}_{i=1}A_{i}$ is any $r$-coloring, there are two sequences $\langle x_{n}\rangle^{\infty}_{n=1}$, and $\langle y_{n}\rangle^{\infty}_{n=1}$ in $\mathbb{N}$, and $i\in\{1,\dots,r\}$ such that $MT(f,\langle x_{n}\rangle^{\infty}_{n=1})\bigcup MT(g,\langle y_{n}\rangle^{\infty}_{n=1})\subseteq A_{i}$.
Proof.
Let $\alpha=\frac{p}{q}$. Let $f,g$ be two polynomials defined by:
$f(x_{1},\dots,x_{m})=\sum a_{i_{1}\cdots i_{m}}x_{1}^{i_{1}}\cdots x_{m}^{i_{m}},$
and
$g(x_{1},\dots,x_{m})=\sum b_{j_{1}}\cdots b_{j_{m}}x_{1}^{j_{1}}\cdots x_{m}^{j_{m}},$
where $b_{i_{1}\cdots i_{m}}=\alpha^{i_{1}+\cdots+i_{m}}a_{i_{1}\cdots i_{m}}$ for every $(i_{1},\dots,i_{m})\in(\mathbb{N}\cup\{0\})^{m}.$
Let $h_{f}:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be a new polynomial defined by $h_{f}(x_{1},\dots,x_{m})=\sum c_{i_{1}\cdots i_{m}}x^{i_{1}}_{1}\cdots x^{i_{m}}_{m}$, where $c_{i_{1}\cdots i_{m}}=p^{i_{1}+\cdots+i_{m}}a_{i_{1}\cdots i_{m}}$ for each $(i_{1},\dots,i_{m})\in(\mathbb{N}\cup\{0\})^{m}$. From Theorem 1.26, there exists $i\in\{1,\dots,r\}$, and a sequence $\langle z_{n}\rangle^{\infty}_{n=1}$ in $\mathbb{N}$ s.t. $MT(h_{f},\langle z_{n}\rangle^{\infty}_{n=1})\subseteq A_{i}$. For each $n\in\mathbb{N}$, define $x_{n}=p\cdot z_{n}$, and $y_{n}=q\cdot z_{n}$. Then, $MT(f,\langle x_{n}\rangle^{\infty}_{n=1})=MT(g,\langle y_{n}\rangle^{\infty}_{n=1})=MT(h_{f},\langle z_{n}\rangle^{\infty}_{n=1})$. This completes the proof. ∎
Acknowledgement
The first named author is supported by NBHM postdoctoral fellowship with reference no: 0204/27/(27)/2023/R & D-II/11927.